Solve the following equations using Wronskian’s method

x²d²y/dx²-3xdy/dx+40y=x²sin(logx)


1
Expert's answer
2021-02-03T01:26:49-0500

Consider Euler equation:

"x^2y''-3xy'+40y=0"

This is Euler Equation.

Let "y(x)=x^r" ; then:

"r(r-1)-3r+40=0"

"r^2-4r+40=0"

"r_{1,2}=\frac{4-\sqrt{16-160}}{2}=2\pm6i"

The general solution is

"y(x)=y_1(x)+y_2(x)=x^2(c_1cos(6lnx)+c_2sin(6lnx))"


Now consider equation

"y''-\frac{3}{x}y+\frac{40}{x^2}=sin(lnx)"

This is equation of the form:

"y''+a(x)y'+b(x)y=f(x)"

The Wronskian is

"W(x)=c\cdot exp\{-\intop a(x)dx\}"

"W(x)=ce^{-3lnx}=c_3x"


The particular solution:

"y_p(x)=-y_1(x)\intop \frac{y_2(x)f(x)}{W(x)}dx+y_2(x)\intop \frac{y_1(x)f(x)}{W(x)}dx"


"\intop \frac{y_2(x)f(x)}{W(x)}dx=\frac{c_2}{c_3}\intop\frac{x^2sin(6lnx)sin(lnx)}{x}dx="

"=\frac{c_2}{c_3}\frac{x^2(265sin(5lnx)-203sin(7lnx)+106cos(5lnx)-58cos(7lnx))}{3074}"


"\intop \frac{y_1(x)f(x)}{W(x)}dx=\frac{c_1}{c_3}\intop\frac{x^2cos(6lnx)sin(lnx)}{x}dx="

"=\frac{c_1}{c_3}\frac{x^2(-106sin(5lnx)+58sin(7lnx)+265cos(5lnx)-203cos(7lnx))}{3074}"


The general solution of the initial equation:

"y(x)=y_1(x)+y_2(x)+y_p(x)"